# Semi-supervised Graph Anomaly Detection via Robust Homophily Learning

### v1

@@ 4.3 GNA: Graph Normality Alignment
where sim($\cdot$,$\cdot$) denotes cosine similarity, and $\tau$ is a temperature hyperparameter.
@@ 5 Experiments
We evaluate RHO on eight real-world GAD datasets of diverse size from different domains, including social networks Reddit [17] and Questions [33], co-review network Amazon [9], co-purchase network Photo [27], collaboration network Tolokers [27], financial networks T-Finance [39], Elliptic [41], and DGraph [14]. See App. B. for more details about the datasets.
@@
The competing methods can be categorized into reconstruction methods (including DOMINANT [6] and AnomalyDAE [11]), one-class classification OCGNN [40], adversarial learning (including AEGIS [5] and GAAN [4]), and affinity maximization method TAM [34], and generative method GGAD [36]. Except for GGAD that is specifically designed for the semi-supervised setting, the other methods are originally unsupervised GAD approaches, which are adapted to the semi-supervised scenario by refining the training set to the labeled normal nodes, following [36]. We also compare RHO against a conventional GCN filter and an advanced spectral-based filter used in BWGNN [39]. Note that although they are fully supervised methods, we utilize only their proposed filter as the encoder and apply it within our semi-supervised setting by optimizing the encoder using the one-class loss. See App. C. for more details about the competing methods. Note that some unsupervised methods, such as CoLA [22], SL-GAD [44], and GRADATE [10], are designed with proxy tasks and cannot be adapted to our setting, which are excluded from our comparison.
@@
Following previous studies [36, 21, 18], we evaluate the detectors using two popular metrics: AUROC (Area Under the Receiver Operating Characteristic curve) and AUPRC (Area Under the Precision-Recall Curve). AUROCassesses the model’s overall ability to distinguish between normal and anomalous nodes across varying thresholds. AUPRC measures the precision-recall tradeoff. For each method, we report the average results over five independent runs.
@@
Our RHO model is implemented in Pytorch 2.0.0 with Python 3.8 and all the experiments are run on GeForce RTX 3090 (24GB). RHO is optimized using Adam [15] optimizer with a defautl learning rate of $35e-$ 3 for all datasets except Elliptic and DGraph that have numerous nodes and sparse edges, in which a smaller learning rate $51e-$ 5 is used for finer optimization. The hyperparameter $\alpha$ is set to $1.0$ for all datasets except the small datasets Reddit and Photo, which require less regularization. It is set to $0.1$ in these two datasets. A detailed analysis on this hyperparameter is presented in Sec. 5.4. Following previous work [36], we randomly sample $R$of the normal nodes as labeled data for training, where $R\in\{5,10,15,20\}$.
@@ 5.1 Main Comparison Results
The main comparison results are shown in Table 1, where all models use 15% labeled normal nodes during training. RHO outperforms the semi-supervised methods on six datasets having maximally 12.19% AUROC and 30.68% AUPRC improvement over the best competing method GGAD. GGAD achieves the highest AUROC on Amazon and Reddit; however, it underperforms RHO on the other datasets. This suggests that generative approaches may not generalize well across different datasets where normal nodes have varying levels of homophily. In contrast, RHO achieves SotA performance on a variety of graph datasets without relying on any generated anomalies, by learning the varying normal patterns from different datasets. The reconstruction-based methods, AnomalyDAE, yields good performance on Tolokers and Amazon, but it still underperforms RHO on most datasets. The main reason is that reconstruction-based methods are prone to overfitting part of the prevalent homophily patterns (e.g., low-frequency nodes) and struggle to handle hard normal samples (e.g., those that have high homophily). In contrast, RHO leverages AdaFreq to effectively learn robust normal patterns from the normal nodes with diverse homophily, leading to substantially improved performance. As for one-class methods, OCGNN is based on non-adaptive filters in learning the normal patterns, it underperforms RHO across all datasets. Moreover, RHO also consistently outperforms GCN and BWGNN under the semi-supervised setting due to its more superior capability in learning heterogeneous frequency components.
@@ 5.3 Ablation Study
Figure 5: AUROC and AUPRC results of RHO w.r.t. hyperparamer $\alpha$
@@ 5.4 Hyperparameter Sensitivity Analysis
We evaluate the sensitivity of RHO w.r.t the hyperparameter $\alpha$ and the training size $R$. More detailed experiments on more datasets and time complexity analysis can be found in App. D.
Figure 6: AUROC w.r.t. data size $R$
Performance w.r.t. Hyperparameter $\alpha$. As shown in Fig. 5, with increasing $\alpha$, the performance on Elliptic, Tolokers, Amazon, Question and T-Finance has different degrees of improvement, suggesting that a stronger consistency constraint between cross-channel and channel-wise representations is beneficial. In contrast, on Reddit and Photo, the performance slightly declines as $\alpha$ increases. This is primarily because these are small datasets with high average feature similarity among the nodes. The representations learned from the two views are already highly similar, and imposing excessive regularization through the normality alignment may lead to overly smoothing features, which may lead to negative effects.
Performance w.r.t. Training Size $R$. To explore the impact of training size $R$, we compare RHO with GGAD, OCGNN, and BWGNN, using varying numbers of training normal nodes, with the results reported in Fig. 6. As the number of the labeled normal nodes increases, RHO generally improves across the datasets, which is consistent with other semi-supervised methods. Under different label rates, our RHO consistently maintains the best performance, showing the superiority of our method.
@@ 6 Conclusion and Future Work
In this paper, we propose RHO, the very first GAD approach designed to learn heterogeneous normal patterns on a set of labeled normal nodes. RHO is implemented by two novel modules, AdaFreq and GNA. AdaFreq learns a set of adaptive spectral filters in both the cross-channel and channel-wise view of node attribute to capture the heterogeneous normal patterns from the given limited labeled normal nodes, while GNA is designed to to enforce the consistency of the learned normal patterns, thereby facilitating the learning of robust normal representations on datasets with different levels of homophily in the normal nodes. This robustness is comprehensively verified by results on eight GAD datasets. A potential limitation of RHO is that it is a transductive method and thus cannot be directly applied in inductive settings. This limitation is shared with most existing GAD methods and is left for future exploration.
@@ Appendix A Theoretical Analysis
A graph filter operation on a graph signal $\mathbf{x}\in\mathbb{R}^{N}$ can be defined as $\mathbf{z}=\mathbf{U}g(\bm{\Lambda})\mathbf{U}^{\top}\mathbf{x}$, where $\mathbf{U}=(\mathbf{u}_{0},\mathbf{u}_{1},\ldots,\mathbf{u}_{N-1})$ is a complete set of orthonormal eigenvectors of the Laplacian matrix, $g(\bm{\Lambda})$ is an adaptive filter with a diagonal matrix and $g(\lambda_{0}),g(\lambda_{1}),\cdots,g(\lambda_{N-1})$ is its main diagonal. In graph signal processing (GSP), $\{\lambda_{i}\}$ and $\{\mathbf{u}_{i}\}$ are called frequencies and frequency components of graph $\mathcal{G}$. Let $\mathcal{V}_{l}$ be the labeled normal nodes in our training data and ${c}$ be a center of a one-class classifier, then the objective function of the one-class classification under semi-supervised graph anomaly detection (GAD) can be defined as:
@@
$\mathcal{L}=\sum_{i\in\mathcal{V}_{l}}||(\mathbf{U}g(\bm{\Lambda})\bm{\beta})_{i}-{c}||^{2}=\sum_{i\in\mathcal{V}_{l}}||{z}_{i}-{c}||^{2},$ (12)
@@
Let $\{\lambda_{m}\}$ and $\{\mathbf{u}_{m}\}$ be the graph frequencies and frequency components respectively, $\beta_{m}$ is the projection coefficient of signal $\mathbf{x}$ onto the $m$-th eigenvector $\mathbf{u}_{m}$, assuming $\beta_{m}>0$, then $g\left(\lambda_{m}\right)=\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\beta_{m}\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}$ holds for the filter $g\left(\lambda\right)$, indicating that frequencies where normal nodes show coherent spectral behavior (i.e., $u_{m}(i)$ values agree in sign/magnitude) are amplified, while inconsistent frequency components are suppressed.
@@
Solving for $g(\lambda_{m})$ yields the closed-form by assuming the projection coefficient $\beta_{m}>0$ and the center $c=1$:
@@
In the theorem, the denominator $\beta_{m}\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}$ represents the weighted total energy of the $m$-th frequency component distributed over the normal nodes. It serves as a normalization factor, reflecting how uniformly the $m$-th frequency component is distributed across the normal nodes. When $u_{m}$ on all labeled normal nodes are in the same direction (same sign), i.e., the frequency component varies consistently across the normal nodes, then $\sum_{i\in\mathcal{V}_{l}}u_{m}(i)\gg 0$, indicating that $g(\lambda_{m})$ generates a strong frequency response to preserve the consistent frequency component across the normal nodes. In contrast, if $u_{m}(i)$ for $i\in\mathcal{V}_{l}$ are highly inconsistent in its direction (different signs), indicating irregular variations among the labeled normal nodes, then $\sum_{i\in\mathcal{V}_{l}}u_{m}(i)\to 0$ and $g(\lambda_{m})$ generates a weak frequency response to restrain the inconsistent frequency component across the normal nodes. Secondly, let $\sigma^{2}$ denote the variance of $u_{m}(i)$ for $i\in\mathcal{V}_{l}$, then if $u_{m}(i)>0$ and $\sum_{i\in\mathcal{V}_{l}}u_{m}(i)$ is fixed, we have $\sigma^{2}\propto-\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}$. The smaller the variance, i.e., the more uniform the magnitude of $u_{m}(i)$), the larger the value of $\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}$, and $g(\lambda_{m})$ will generate a strong frequency response. As a result, frequency components where normal nodes exhibit coherent spectral behavior—where the values of $u_{m}(i)$ are consistent in sign and magnitude—are amplified, whereas those with inconsistent patterns are suppressed.
@@ Appendix B Detailed Description of Datasets
  • Reddit [17]: It is a user-subreddit graph which captures one month’s worth of posts shared across various subreddits at Reddit. The node represents the users, and the text of each post is transformed into a feature vector and the features of the user and subreddits are the feature summation of the post they have posted. The anomalies are the used who have been banned by the platform.
  • Tolokers [27]: It is obtained from the Toloka crowdsourcing platform, where the node represents the person who has participated in the selected project, and the edge represents two workers work on the same task. The attributes of the node are the profile and task performance statistics of workers.
  • Photo [27]: It is obtained from an Amazon co-purchase network where the node represents the product and the edge represents the co-purchase relationship. The bag-of-words representation of the user’s comments is used as the attribute of the node.
  • Amazon [9]: It is a co-review network obtained from the Musical Instrument category on Amazon.com. There are also three relations: U-P-U (users reviewing at least one same product), U-S-U (users having at least one same star rating within one week), and U-V-U (users with top-5% mutual review similarities).
  • Elliptic [41]: It is a Bitcoin transaction network in which each node represents a transaction and an edge indicates a flow of Bitcoin currency. The Bitcoin transaction is mapped to real-world entities associated with licit categories in this dataset.
  • Question [33]: It is a question answering network which is collected from the website Yandex Q where the node represents the user and the edge connecting the node represents the user who has answered other’s questions during a one-year period. The attribute of nodes is the mean of FastText embeddings for words in the description of the user. For the user without a description, the additional binary features are employed as the feature of the user.
  • T-Finance [39]: It is a financial transaction network where the node represents an anonymous account and the edge represents two accounts that have transaction records. Some attributes of logging like registration days, logging activities, and interaction frequency, etc, are used as the features of each account. Users are labeled as anomalies if they fall into categories such as fraud, money laundering, or online gambling.
  • DGraph [14]: It is a large-scale attributed graph with millions of nodes and edges where the node represents a user account in a financial company and the edge represents that the user was added to another account as an emergency contact. The feature of each node represents the user’s profile information, including age, gender, and other demographic attributes. Users with a history of overdue payments are labeled as anomalies.

### v2

@@ 4.3 GNA: Graph Normality Alignment
where sim($\cdot$$\cdot$) denotes cosine similarity, and $\tau$ is a temperature hyperparameter.
@@ 4.4 Training and Inference
4.5 Complexity Analysis
The computational complexity of RHO consists of three parts: (1) The two-layer MLP used for initial feature transformation has a complexity of $\mathcal{O}(Nd(M+d))$, where $N$ denotes the number of nodes, $M$ is the input feature dimension, and $d$ represents the hidden dimension. (2) RHO employs two independent views, including the cross-channel and channel-wise views. Each view incurs a computational cost of $\mathcal{O}(|\mathcal{E}|d+Nd^{2})$, where $|\mathcal{E}|$ is the number of edges. The overall complexity of this component remains $\mathcal{O}(2*(|\mathcal{E}|d+Nd^{2}))$ in practice. (3) In GNA, the projection head has a complexity of $\mathcal{O}(Nd^{2})$, and the alignment loss computation has a complexity of $\mathcal{O}(Nbd)$, where $b$ denotes the batch size. Therefore, the overall time complexity is $\mathcal{O}(NMd+2|\mathcal{E}|d+4Nd^{2}+Nbd))$. Empirical results for running time can be found in App. D.4.
@@ 5 Experiments
We evaluate RHO on eight real-world GAD datasets of diverse size from different domains, including social networks Reddit [kumar2019predicting] and Questions [platonov2023critical], co-review network Amazon [dou2020enhancing], co-purchase network Photo [mcauley2015image], collaboration network Tolokers [mcauley2015image], financial networks T-Finance [tang2022rethinking], Elliptic [weber2019anti], and DGraph [huang2022DGraph]. See App. B for more details about the datasets.
@@
The competing methods can be categorized into reconstruction methods (including DOMINANT [ding2019deep] and AnomalyDAE [fan2020anomalydae]), one-class classification OCGNN [wang2021one], adversarial learning (including AEGIS [ding2021inductive] and GAAN [chen2020generative]), affinity maximization method TAM [hezhe2023truncated], generative method GGAD [qiao2024generative], partitioning message passing method (PMP) [zhuo2024partitioning], and data augmentation method CONSISGAD [chen2024consistency]. Except for GGAD that is specifically designed for the semi-supervised setting, the other methods are originally unsupervised or fully supervised GAD approaches, which are adapted to the semi-supervised scenario by refining the training set to the labeled normal nodes, following [qiao2024generative]. We also compare RHO against a conventional GCN filter and an advanced spectral-based filter used in BWGNN [tang2022rethinking]. Note that for the fully supervised methods, we utilize only their proposed filter as the encoder and apply it within our semi-supervised setting by optimizing the encoder using the one-class loss. See App. C for more details about the competing methods. Note that some unsupervised methods, such as CoLA [liu2021anomaly], SL-GAD [zheng2021generative], and GRADATE [duan2023graph], are designed with proxy tasks and cannot be adapted to our setting, which are excluded from our comparison.
@@
Following previous studies [qiao2024generative, liu2024arc, liu2023towards], we evaluate the detectors using two popular metrics: AUROC (Area Under the Receiver Operating Characteristic curve) and AUPRC (Area Under the Precision-Recall Curve). AUROC assesses the model’s overall ability to distinguish between normal and anomalous nodes across varying thresholds. AUPRC measures the precision-recall tradeoff. For each method, we report the average results over five independent runs.
@@
The RHO model is implemented in PyTorch 2.0.0 with Python 3.8 and executed on GeForce RTX 3090 GPU (24 GB). RHO is trained using the Adam optimizer [kinga2015method] with a weight decay of $5e^{-5}$. We set the default learning rate to $5e^{-3}$. Nevertheless, owing to the variations in edge density across different graphs, we find that models trained on sparser graphs are generally more sensitive to large learning rates and thus benefit from smaller ones to ensure stable convergence. Therefore, we use a learning rate of $5e^{-4}$ for the Elliptic and Question datasets, and further decreased to $5e^{-6}$ for the extremely sparse dataset, DGraph. The hyperparameter $\alpha$ is set to $1.0$ for all datasets except the small datasets Reddit and Photo, which require less regularization. It is set to $0.1$ in these two datasets. A detailed analysis of this hyperparameter is presented in Sec. 5.4. Following previous work [qiao2024generative], we randomly sample RR% of the normal nodes as labeled data for training, where $R\in\{5,10,15,20\}$. To ensure a fair comparison, we obtain the publicly-available official source code of all competitors and execute these models using the parameter settings suggested by their authors.
@@ 5.1 Main Comparison Results
The main comparison results are shown in Table 1, where all models use 15% labeled normal nodes during training. RHO outperforms the semi-supervised methods on six datasets having maximally 12.19% AUROC and 30.68% AUPRC improvement over the best competing method GGAD. GGAD achieves the highest AUROC on Amazon and Reddit; however, it underperforms RHO on the other datasets. This suggests that generative approaches may not generalize well across different datasets where normal nodes have varying levels of homophily. In contrast, RHO achieves SotA performance on a variety of graph datasets without relying on any generated anomalies, by learning the varying normal patterns from different datasets. The fully supervised methods PMP and CONSISGAD perform less effectively when applied in the semi-supervised setting, since the lack of labeled anomalies prevents them from leveraging crucial abnormality information. The reconstruction-based method, AnomalyDAE, yields good performance on Tolokers and Amazon, but it still underperforms RHO on most datasets. The main reason is that reconstruction-based methods are prone to overfitting part of the prevalent homophily patterns (e.g., low-frequency nodes) and struggle to handle hard normal samples (e.g., those that have high homophily). In contrast, RHO leverages AdaFreq to effectively learn robust normal patterns from the normal nodes with diverse homophily, leading to substantially improved performance. As for one-class methods, OCGNN is based on non-adaptive filters in learning the normal patterns; it underperforms RHO across all datasets. Moreover, RHO also consistently outperforms GCN and BWGNN under the semi-supervised setting due to its superior capability in learning heterogeneous frequency components.
@@ 5.4 Hyperparameter Sensitivity Analysis
We evaluate the sensitivity of RHO w.r.t. the hyperparameter $\alpha$ and the training size $R$. Detailed results on more datasets can be found in App. D.
Figure 5: AUROC and AUPRC results of RHO w.r.t. hyperparamer $\alpha$.
Performance w.r.t. Hyperparameter $\alpha$.
As shown in Fig. 5, with increasing $\alpha$, the performance on Elliptic, Tolokers, Amazon, Question and T-Finance has different degrees of improvement, suggesting that a stronger consistency constraint between cross-channel and channel-wise representations is beneficial. In contrast, on Reddit and Photo, the performance slightly declines as $\alpha$ increases. This is primarily because these are small datasets with high average feature similarity among the nodes. The representations learned from the two views are already highly similar, and imposing excessive regularization through the normality alignment may lead to overly smoothing features, which may lead to negative effects.
Figure 6: AUROC w.r.t. data size $R$.
Performance w.r.t. Training Size $R$.
To explore the impact of training size $R$, we compare RHO with GGAD, OCGNN, and BWGNN, using varying numbers of training normal nodes, with the results reported in Fig. 6. As the number of the labeled normal nodes increases, RHO generally improves across the datasets, which is consistent with other semi-supervised methods. Notably, RHO demonstrates a stable and remarkable improvement trend on the Elliptic dataset, even when the amount of labeled data is limited. This observation indicates that our model can effectively exploit the normality patterns under sparse supervision. Under different label rates, our RHO consistently maintains the best performance, showing the superiority of our method.
@@ 6 Conclusion and Future Work
In this paper, we propose RHO, the very first GAD approach designed to learn heterogeneous normal patterns on a set of labeled normal nodes. RHO is implemented by two novel modules, AdaFreq and GNA. AdaFreq learns a set of adaptive spectral filters in both the cross-channel and channel-wise view of node attribute to capture the heterogeneous normal patterns from the given limited labeled normal nodes, while GNA is designed to enforce the consistency of the learned normal patterns, thereby facilitating the learning of robust normal representations on datasets with different levels of homophily in the normal nodes. This robustness is comprehensively verified by results on eight GAD datasets. A potential limitation of RHO is that it is a transductive method and thus cannot be directly applied in inductive settings. This limitation is shared with most existing GAD methods and is left for future exploration.
Acknowledgments
This research is supported by the Ministry of Education, Singapore under its Tier-1 Academic Research Fund (24-SIS-SMU-008), A*STAR under its MTC YIRG Grant (M24N8c0103), and the Lee Kong Chian Fellowship.
@@ Appendix A Theoretical Analysis
A graph filter operation on a graph signal $\mathbf{x}\in\mathbb{R}^{N}$ can be defined as $\mathbf{z}=\mathbf{U}g(\bm{\Lambda})\mathbf{U}^{\top}\mathbf{x}$, where $\mathbf{U}=(\mathbf{u}_{0},\mathbf{u}_{1},\ldots,\mathbf{u}_{N-1})$ is a complete set of orthonormal eigenvectors of the Laplacian matrix, $g(\bm{\Lambda})$ is an adaptive filter with a diagonal matrix and $g(\lambda_{0}),g(\lambda_{1}),\cdots,g(\lambda_{N-1})$ is its main diagonal. In graph signal processing (GSP), $\{\lambda_{i}\}$ and $\{\mathbf{u}_{i}\}$ are called frequencies and frequency components of graph $\mathcal{G}$. Let $\mathcal{V}_{l}$ be the labeled normal nodes in our training data and c{c} be a center of a one-class classifier, then the objective function of the one-class classification under semi-supervised graph anomaly detection (GAD) can be defined as:
@@
[…]
@@
Let $\{\lambda_{m}\}$ and $\{\mathbf{u}_{m}\}$ be the graph frequencies and frequency components respectively, $\beta_{m}$ is the projection coefficient of signal $\mathbf{x}$ onto the $m$-th eigenvector $\mathbf{u}_{m}$, then $g\left(\lambda_{m}\right)=\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\beta_{m}\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}$ holds for the filter $g\left(\lambda\right)$, indicating that frequencies where normal nodes show coherent spectral behavior (i.e., $u_{m}(i)$ values agree in sign/magnitude) are amplified, while inconsistent frequency components are suppressed.
@@
Solving for $g(\lambda_{m})$ yields the closed-form by assuming the center $c=1$:
@@
In the theorem, the denominator $\beta_{m}\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}$ represents the weighted total energy of the $m$-th frequency component distributed over the normal nodes. It serves as a normalization factor, reflecting how uniformly the $m$-th frequency component is distributed across the normal nodes. When the $m$-th frequency components exhibit high consistency across the labeled normal nodes, e.g., $u_{m}(i)$ have the same sign (all positive or all negative) and similar amplitude for $i\in\mathcal{V}_{l}$, then the ratio $\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\beta_{m}\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}$ will be either much greater than 0 (case 1: $\beta_{m}>0$ and $\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}\gg 0$, case 2: $\beta_{m}<0$ and $\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}\ll 0$) or much less than 0 (case 1: $\beta_{m}>0$ and $\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}\ll 0$, case 2: $\beta_{m}<0$ and $\frac{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)}{\sum_{i\in\mathcal{V}_{l}}u_{m}(i)^{2}}\gg 0$), resulting in a frequency positively or negatively enhancing response $g(\lambda_{m})\gg 0$ or $g(\lambda_{m})\ll 0$. Only when the frequency components are highly inconsistent, e.g., $u_{m}(i)$ exhibits opposite signs across the labeled normal nodes, these inconsistent components will cancel each other out (i.e., $\sum_{i\in\mathcal{V}_{l}}u_{m}(i)\rightarrow 0$), resulting in a frequency suppressing response $g(\lambda_{m})\to 0$. To summarize, minimizing the one-class loss enables an adaptive filtering mechanism to automatically enhance frequency components that are consistent across normal nodes while suppressing the inconsistent components.
@@ Appendix B Detailed Description of Datasets
  • Reddit [kumar2019predicting]: It is a user-subreddit graph which captures one month’s worth of posts shared across various subreddits at Reddit. The node represents the users, and the text of each post is transformed into a feature vector and the features of the user and subreddits are the feature summation of the post they have posted. The anomalies are the used who have been banned by the platform.
  • Tolokers [mcauley2015image]: It is obtained from the Toloka crowdsourcing platform, where the node represents the person who has participated in the selected project, and the edge represents two workers work on the same task. The attributes of the node are the profile and task performance statistics of workers.
  • Photo [mcauley2015image]: It is obtained from an Amazon co-purchase network where the node represents the product and the edge represents the co-purchase relationship. The bag-of-words representation of the user’s comments is used as the attribute of the node.
  • Amazon [dou2020enhancing]: It is a co-review network obtained from the Musical Instrument category on Amazon.com. There are also three relations: U-P-U (users reviewing at least one same product), U-S-U (users having at least one same star rating within one week), and U-V-U (users with top-5% mutual review similarities).
  • Elliptic [weber2019anti]: It is a Bitcoin transaction network in which each node represents a transaction and an edge indicates a flow of Bitcoin currency. The Bitcoin transaction is mapped to real-world entities associated with licit categories in this dataset.
  • Question [platonov2023critical]: It is a question answering network which is collected from the website Yandex Q where the node represents the user and the edge connecting the node represents the user who has answered other’s questions during a one-year period. The attribute of nodes is the mean of FastText embeddings for words in the description of the user. For the user without a description, the additional binary features are employed as the feature of the user.
  • T-Finance [tang2022rethinking]: It is a financial transaction network where the node represents an anonymous account and the edge represents two accounts that have transaction records. Some attributes of logging like registration days, logging activities, and interaction frequency, etc, are used as the features of each account. Users are labeled as anomalies if they fall into categories such as fraud, money laundering, or online gambling.
  • DGraph [huang2022DGraph]: It is a large-scale attributed graph with millions of nodes and edges where the node represents a user account in a financial company and the edge represents that the user was added to another account as an emergency contact. The feature of each node represents the user’s profile information, including age, gender, and other demographic attributes. Users with a history of overdue payments are labeled as anomalies.